# Optimal frames for Phase Retrieval from Edge Vectors of Optimal Polygons

### v1

@@ 2.3. Basic properties of convex polygons
[26] The width function of a convex polygon ${P}\subset\mathbb{R}^{2}$ in the direction $\bm{u}\in\mathbb{S}^{1}$ is defined by $w({P},\bm{u}):=h({P},\bm{u})+h({P},-\bm{u})$, where $h({P},\bm{z}):=\max_{\bm{x}\in{P}}\langle\bm{x},\bm{z}\rangle$ for any $\bm{z}\in\mathbb{S}^{1}$.
@@
[26] For any convex polygon ${P}\subset\mathbb{R}^{2}$, we have $\operatorname{diam}({P})=\max_{\bm{u}\in\mathbb{S}^{1}}w({P},\bm{u}).$
@@
Assume that $E=\{\bm{e}_{1},\dotsc,\bm{e}_{m}\}\subset\mathbb{R}^{2}$ is the edge set of a convex $m$-gon ${P}\subset\mathbb{R}^{2}$. Then
$r({P})=\frac{\underset{\bm{u}\in\mathbb{S}^{1}}{\max}\,\sum_{j=1}^{m}{|\langle\bm{e}_{j},\bm{u}\rangle|}}{2\sum_{j=1}^{m}\|\bm{e}_{j}\|_{2}}.$ (2.7)
@@
We first consider the case when $P$ is strongly convex. Since $r(P)=\frac{\mathrm{diam}(P)}{\mathrm{perim}(P)}$ and the perimeter $\mathrm{perim}(P)$ equals $\sum_{j=1}^{m}\|\bm{e}_{j}\|_{2}$, it is enough to prove
$\operatorname{diam}(P)=\frac{1}{2}\max_{\bm{u}\in\mathbb{S}^{1}}\,\sum_{j=1}^{m}{|\langle\bm{e}_{j},\bm{u}\rangle|}.$ (2.8)
Without loss of generality, we assume that $\bm{e}_{1},\dotsc,\bm{e}_{m}$ are listed in counterclockwise order. Let $\bm{v}_{1},\dotsc,\bm{v}_{m}$ be the vertices of $P$ such that $\bm{e}_{j}=\bm{v}_{j+1}-\bm{v}_{j}$ for each $j\in[m]$, where we denote $\bm{v}_{m+1}=\bm{v}_{1}$. For any $\bm{u}\in\mathbb{S}^{1}$, since $P$ is convex, there exist ${j_{1}},\,j_{2}\in[m]$ depending on $\bm{u}$ such that
$\bm{v}_{j_{1}}\in\operatorname*{argmin}_{\bm{x}\in P}\langle\bm{x},\bm{u}\rangle\quad\text{and}\quad\bm{v}_{j_{2}}\in\operatorname*{argmax}_{\bm{x}\in P}\langle\bm{x},\bm{u}\rangle.$ (2.9)
@@
$\operatorname{diam}(P)=\max_{\bm{u}\in\mathbb{S}^{1}}w(P,\bm{u})$ $=\max_{\bm{u}\in\mathbb{S}^{1}}(h(P,\bm{u})+h(P,-\bm{u}))=\max_{\bm{u}\in\mathbb{S}^{1}}(\max_{\bm{x}\in{P}}\langle\bm{x},\bm{u}\rangle+\max_{\bm{x}\in{P}}\langle\bm{x},-\bm{u}\rangle)$ (2.10)
$=\max_{\bm{u}\in\mathbb{S}^{1}}(\max_{\bm{x}\in{P}}\langle\bm{x},\bm{u}\rangle-\min_{\bm{x}\in{P}}\langle\bm{x},\bm{u}\rangle)=\max_{\bm{u}\in\mathbb{S}^{1}}\,{\langle\bm{v}_{j_{2}}-\bm{v}_{j_{1}},\bm{u}\rangle}.$
@@
$\langle\bm{e}_{j},\bm{u}\rangle\begin{cases}\geq 0&\text{if $j_{1}\leq j\leq j_{2}-1$,}\\ \leq 0&\text{else.}\end{cases}$ (2.11)
@@
$\begin{split}\langle\bm{v}_{j_{2}}-\bm{v}_{j_{1}},\bm{u}\rangle=\langle\sum_{j=j_{1}}^{j_{2}-1}{\bm{e}_{j}},\bm{u}\rangle\overset{(a)}{=}\langle\frac{1}{2}\sum_{j=j_{1}}^{j_{2}-1}{\bm{e}_{j}}-\frac{1}{2}\sum_{j=1}^{j_{1}-1}{\bm{e}_{j}}-\frac{1}{2}\sum_{j=j_{2}}^{m}{\bm{e}_{j}},\bm{u}\rangle\overset{(b)}{=}\frac{1}{2}\sum_{j=1}^{m}{|\langle\bm{e}_{j},\bm{u}\rangle|},\end{split}$ (2.12)
where $(a)$ follows from $\sum_{j=1}^{m}\bm{e}_{j}=\bm{0}$ and $(b)$ follows from (2.11). Substituting (2.12) into (2.10), we arrive at (2.8).
It remains to prove (2.11). We first prove $\langle\bm{e}_{j},\bm{u}\rangle\geq 0$ for each $j_{1}\leq j\leq j_{2}-1$. By the definition of $j_{1}$ and $j_{2}$, we have
$\langle\bm{e}_{j_{1}},\bm{u}\rangle=\langle\bm{v}_{j_{1}+1},\bm{u}\rangle-\langle\bm{v}_{j_{1}},\bm{u}\rangle\geq 0\quad\text{and}\quad\langle\bm{e}_{j_{2}-1},\bm{u}\rangle=\langle\bm{v}_{j_{2}},\bm{u}\rangle-\langle\bm{v}_{j_{2}-1},\bm{u}\rangle\geq 0.$ (2.13)
Suppose, for contradiction, that there exists an integer $s$ with $j_{1}+1\leq s\leq j_{2}-2$ such that $\langle\bm{e}_{s},\bm{u}\rangle<0$. Then there must exist an integer $t\in\{s+1,\dotsc,j_{2}-1\}$ satisfying
$\langle\bm{e}_{t-1},\bm{u}\rangle<0,\quad\langle\bm{e}_{t},\bm{u}\rangle\geq 0,$ (2.14)
since otherwise we have $\langle\bm{e}_{i},\bm{u}\rangle<0$ for each $s+1\leq i\leq j_{2}-1$, which contradicts with (2.13). Since $P$ is strongly convex, for any $\bm{x}\in P$ there exist $\gamma_{1},\gamma_{2}\geq 0$ depending on $\bm{x}$ such that
$\bm{x}-\bm{v}_{t}=\gamma_{1}\cdot(-\bm{e}_{t-1})+\gamma_{2}\cdot\bm{e}_{t}.$
Taking inner products with $\bm{u}$ and using (2.14), we have
$\langle\bm{x},\bm{u}\rangle-\langle\bm{v}_{t},\bm{u}\rangle=\gamma_{1}\langle-\bm{e}_{t-1},\bm{u}\rangle+\gamma_{2}\langle\bm{e}_{t},\bm{u}\rangle\geq 0,\quad\forall\bm{x}\in P,$
@@
$\langle\bm{v}_{j_{1}},\bm{u}\rangle=\langle\bm{v}_{t},\bm{u}\rangle={\min}_{\bm{x}\in P}\langle\bm{x},\bm{u}\rangle.$ (2.15)
It is well known that a linear function on a convex polygon attains its extrema at either a vertex or along an edge of the polygon [27]. Hence, we must have $t=j_{1}+1$. This contradicts with $t\geq s+1\geq j_{1}+2$. Therefore, we have $\langle\bm{e}_{j},\bm{u}\rangle\geq 0$ for each $j_{1}\leq j\leq j_{2}-1$. Similarly, we can prove that $\langle\bm{e}_{j},\bm{u}\rangle\leq 0$ for each $j<j_{1}$ and $j\geq j_{2}$. Hence, we arrive at (2.11).
We next consider the case when $P$ is not strictly convex. By replacing all vectors in the same direction with their sum, we obtain a strictly convex $k$-gon $\widehat{P}\in\mathcal{P}_{k}$ with the edge set $\widehat{E}=\{\widehat{\bm{e}}_{1},\dotsc,\widehat{\bm{e}}_{k}\}$, where $k$ is an integer less than $m$. Note that $\sum_{j=1}^{m}\|\bm{e}_{j}\|_{2}=\sum_{j=1}^{k}\|\widehat{\bm{e}}_{j}\|_{2}$ and $\sum_{j=1}^{m}|\langle\bm{e}_{j},\bm{u}\rangle|=\sum_{j=1}^{k}|\langle\widehat{\bm{e}}_{j},\bm{u}\rangle|$ for any $\bm{u}\in\mathbb{S}^{1}$. Hence, we have
$r(P)=r(\widehat{P})=\frac{\underset{\bm{u}\in\mathbb{S}^{1}}{\max}\,\sum_{j=1}^{m}{|\langle\widehat{\bm{e}}_{j},\bm{u}\rangle|}}{2\sum_{j=1}^{m}\|\widehat{\bm{e}}_{j}\|_{2}}=\frac{\underset{\bm{u}\in\mathbb{S}^{1}}{\max}\,\sum_{j=1}^{m}{|\langle\bm{e}_{j},\bm{u}\rangle|}}{2\sum_{j=1}^{m}\|\bm{e}_{j}\|_{2}}.$
@@
[21] Suppose $\bm{u}_{1},\dotsc,\bm{u}_{m}\in\mathbb{R}^{2}$ are distinct unit vectors that span $\mathbb{R}^{2}$, and suppose that $\alpha_{1},\dotsc,\alpha_{m}>0$. Then there exists a strictly convex $m$-gon P{P} in $\mathbb{R}^{2}$, having edge unit outer normals $\bm{u}_{1},\dotsc,\bm{u}_{m}$ and corresponding edge lengths $\alpha_{1},\dotsc,\alpha_{m}>0$, if and only if $\sum_{j=1}^{m}\alpha_{j}\bm{u}_{j}=\bm{0}.$ Moreover, such a polygon P{P} is unique up to translation.
@@
Suppose $E=\{\bm{e}_{1},\dotsc,\bm{e}_{m}\}\subset\mathbb{R}^{2}$. Then there exists a strictly convex $m$-gon P{P} in $\mathbb{R}^{2}$, having the edge set $E$ if and only if $\bm{e}_{1},\dotsc,\bm{e}_{m}$ are nonzero vectors with pairwise distinct directions and $\sum_{j=1}^{m}\bm{e}_{j}=\bm{0}.$ Moreover, such a polygon P{P} is unique up to translation.
@@
Assume that $\sum_{j=1}^{m}\bm{e}_{j}=\bm{0}$. For all $j\in[m]$, rotating each vector $\frac{\bm{e}_{j}}{\|\bm{e}_{j}\|_{2}}$ counterclockwise by $\frac{\pi}{2}$ yields $\bm{u}_{j}$, while preserving the zero-sum property $\sum_{j=1}^{m}\|\bm{e}_{j}\|_{2}\bm{u}_{j}=\bm{0}$. It is easy to verify that $\bm{u}_{1},\dotsc,\bm{u}_{m}\in\mathbb{R}^{2}$ are distinct unit vectors that span $\mathbb{R}^{2}$. Therefore we can utilize Theorem 2.3 to obtain a strictly convex $m$-gon P{P} with edge unit outer normals $\bm{u}_{1},\dotsc,\bm{u}_{m}$ and corresponding edge lengths $\|\bm{e}_{1}\|_{2},\dotsc,\|\bm{e}_{m}\|_{2}$, i.e., with edge vectors $\bm{e}_{1},\dotsc,\bm{e}_{m}$. Moreover, P{P} is unique, disregarding the translation of polygons.
Conversely, assume that P{P} is a strictly convex $m$-gon with the edge set $E$. Assume that $\bm{e}_{1},\dotsc,\bm{e}_{m}$ are arranged counterclockwise. It is clear that $\sum_{j=1}^{m}\bm{e}_{j}=\bm{0}$ and each $\bm{e}_{i}$ is nonzero. Without loss of generality, it is enough to prove that $\bm{e}_{1}$ and $\bm{e}_{i}$ have distinct directions for any $i>1$. For each $i\in\{2,\ldots,m\}$, define $\alpha_{i}$ as the counterclockwise angle from $\bm{e}_{1}$ to $\bm{e}_{i}$. Since $P$ is strictly convex, we have $\alpha_{i}\in(0,2\pi)$, and $\alpha_{i}$ is strictly increasing as $i$ increases. Thus, there does not exist an integer $i\in\{2,\ldots,m\}$ such that $\alpha_{i}=0$, i.e., $\bm{e}_{1}$ and $\bm{e}_{j}$ have the same direction. This completes the proof.

### v2

@@ 2.3. Basic properties of convex polygons
[26] The width function of a convex polygon ${P}\subset\mathbb{R}^{2}$ in the direction $\boldsymbol{u}\in\mathbb{S}^{1}$ is defined by $w({P},\boldsymbol{u}):=h({P},\boldsymbol{u})+h({P},-\boldsymbol{u})$, where $h({P},\boldsymbol{z}):=\max_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},\boldsymbol{z}\rangle$ for any $\boldsymbol{z}\in\mathbb{S}^{1}$.
@@
[26] For any convex polygon ${P}\subset\mathbb{R}^{2}$, we have
$\operatorname{diam}({P})=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}w({P},\boldsymbol{u}).$
@@
Assume that $E=\{\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}\}\subset\mathbb{R}^{2}$ is the edge set of a convex $m$-gon ${P}\subset\mathbb{R}^{2}$. Then
$r({P})=\frac{\underset{\boldsymbol{u}\in\mathbb{S}^{1}}{\max}\,\sum_{j=1}^{m}{|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|}}{2\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}}.$ (2.7)
@@
We first consider the case when $P$ is strongly convex. Since $r(P)=\frac{\mathrm{diam}(P)}{\mathrm{perim}(P)}$ and the perimeter $\mathrm{perim}(P)$ equals $\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}$, it is enough to prove
$\operatorname{diam}(P)=\frac{1}{2}\max_{\boldsymbol{u}\in\mathbb{S}^{1}}\,\sum_{j=1}^{m}{|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|}.$ (2.8)
Without loss of generality, we assume that $\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}$ are listed in counterclockwise order. Let $\boldsymbol{v}_{1},\dotsc,\boldsymbol{v}_{m}$ be the vertices of $P$ such that $\boldsymbol{e}_{j}=\boldsymbol{v}_{j+1}-\boldsymbol{v}_{j}$ for each $j\in[m]$, where we denote $\boldsymbol{v}_{m+1}=\boldsymbol{v}_{1}$. For any $\boldsymbol{u}\in\mathbb{S}^{1}$, since $P$ is convex, there exist ${j_{1}},\,j_{2}\in[m]$ depending on $\boldsymbol{u}$ such that
$\boldsymbol{v}_{j_{1}}\in\operatorname*{argmin}_{\boldsymbol{x}\in P}\langle\boldsymbol{x},\boldsymbol{u}\rangle\quad\text{and}\quad\boldsymbol{v}_{j_{2}}\in\operatorname*{argmax}_{\boldsymbol{x}\in P}\langle\boldsymbol{x},\boldsymbol{u}\rangle.$ (2.9)
@@
$\operatorname{diam}(P)=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}w(P,\boldsymbol{u})$ $=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}(h(P,\boldsymbol{u})+h(P,-\boldsymbol{u}))$ (2.10)
$=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}(\max_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},\boldsymbol{u}\rangle+\max_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},-\boldsymbol{u}\rangle)$
$=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}(\max_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},\boldsymbol{u}\rangle-\min_{\boldsymbol{x}\in{P}}\langle\boldsymbol{x},\boldsymbol{u}\rangle)$
$=\max_{\boldsymbol{u}\in\mathbb{S}^{1}}\,{\langle\boldsymbol{v}_{j_{2}}-\boldsymbol{v}_{j_{1}},\boldsymbol{u}\rangle}.$
@@
$\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle\begin{cases}\geq 0&\text{if $j_{1}\leq j\leq j_{2}-1$,}\\ \leq 0&\text{else.}\end{cases}$ (2.11)
@@
$\langle\boldsymbol{v}_{j_{2}}-\boldsymbol{v}_{j_{1}},\boldsymbol{u}\rangle=\langle\sum_{j=j_{1}}^{j_{2}-1}{\boldsymbol{e}_{j}},\boldsymbol{u}\rangle$ $\overset{(a)}{=}\langle\frac{1}{2}\sum_{j=j_{1}}^{j_{2}-1}{\boldsymbol{e}_{j}}-\frac{1}{2}\sum_{j=1}^{j_{1}-1}{\boldsymbol{e}_{j}}-\frac{1}{2}\sum_{j=j_{2}}^{m}{\boldsymbol{e}_{j}},\boldsymbol{u}\rangle$ (2.12)
$\overset{(b)}{=}\frac{1}{2}\sum_{j=1}^{m}{|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|},$
where $(a)$ follows from $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$ and $(b)$ follows from (2.11). Substituting (2.12) into (2.10), we arrive at (2.8).
It remains to prove (2.11). We first prove $\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle\geq 0$ for each $j_{1}\leq j\leq j_{2}-1$. By the definition of $j_{1}$ and $j_{2}$, we have
$\langle\boldsymbol{e}_{j_{1}},\boldsymbol{u}\rangle=\langle\boldsymbol{v}_{j_{1}+1},\boldsymbol{u}\rangle-\langle\boldsymbol{v}_{j_{1}},\boldsymbol{u}\rangle\geq 0,\;\,\langle\boldsymbol{e}_{j_{2}-1},\boldsymbol{u}\rangle=\langle\boldsymbol{v}_{j_{2}},\boldsymbol{u}\rangle-\langle\boldsymbol{v}_{j_{2}-1},\boldsymbol{u}\rangle\geq 0.$ (2.13)
Suppose, for contradiction, that there exists an integer $s$ with $j_{1}+1\leq s\leq j_{2}-2$ such that $\langle\boldsymbol{e}_{s},\boldsymbol{u}\rangle<0$. Then there must exist an integer $t\in\{s+1,\dotsc,j_{2}-1\}$ satisfying
$\langle\boldsymbol{e}_{t-1},\boldsymbol{u}\rangle<0,\quad\langle\boldsymbol{e}_{t},\boldsymbol{u}\rangle\geq 0,$ (2.14)
since otherwise we have $\langle\boldsymbol{e}_{i},\boldsymbol{u}\rangle<0$ for each $s+1\leq i\leq j_{2}-1$, which contradicts with (2.13). Since $P$ is strongly convex, for any $\boldsymbol{x}\in P$ there exist $\gamma_{1},\gamma_{2}\geq 0$ depending on $\boldsymbol{x}$ such that
$\boldsymbol{x}-\boldsymbol{v}_{t}=\gamma_{1}\cdot(-\boldsymbol{e}_{t-1})+\gamma_{2}\cdot\boldsymbol{e}_{t}.$
Taking inner products with $\boldsymbol{u}$ and using (2.14), we have
$\langle\boldsymbol{x},\boldsymbol{u}\rangle-\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle=\gamma_{1}\langle-\boldsymbol{e}_{t-1},\boldsymbol{u}\rangle+\gamma_{2}\langle\boldsymbol{e}_{t},\boldsymbol{u}\rangle\geq 0,\quad\forall\boldsymbol{x}\in P,$
@@
$\langle\boldsymbol{v}_{j_{1}},\boldsymbol{u}\rangle=\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle={\min}_{\boldsymbol{x}\in P}\langle\boldsymbol{x},\boldsymbol{u}\rangle.$ (2.15)
It is well known that a linear function on a convex polygon attains its extrema at either a vertex or along an edge of the polygon [27]. Hence, we must have $t=j_{1}+1$. This contradicts with $t\geq s+1\geq j_{1}+2$. Therefore, we have $\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle\geq 0$ for each $j_{1}\leq j\leq j_{2}-1$. Similarly, we can prove that $\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle\leq 0$ for each $j<j_{1}$ and $j\geq j_{2}$. Hence, we arrive at (2.11).
We next consider the case when $P$ is not strictly convex. By replacing all vectors in the same direction with their sum, we obtain a strictly convex $k$-gon $\widehat{P}\in\mathcal{P}_{k}$ with the edge set $\widehat{E}=\{\widehat{\boldsymbol{e}}_{1},\dotsc,\widehat{\boldsymbol{e}}_{k}\}$, where $k$ is an integer less than $m$. Note that $\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}=\sum_{j=1}^{k}\|\widehat{\boldsymbol{e}}_{j}\|_{2}$ and $\sum_{j=1}^{m}|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|=\sum_{j=1}^{k}|\langle\widehat{\boldsymbol{e}}_{j},\boldsymbol{u}\rangle|$ for any $\boldsymbol{u}\in\mathbb{S}^{1}$. Hence, we have
$r(P)=r(\widehat{P})=\frac{\underset{\boldsymbol{u}\in\mathbb{S}^{1}}{\max}\,\sum_{j=1}^{m}{|\langle\widehat{\boldsymbol{e}}_{j},\boldsymbol{u}\rangle|}}{2\sum_{j=1}^{m}\|\widehat{\boldsymbol{e}}_{j}\|_{2}}=\frac{\underset{\boldsymbol{u}\in\mathbb{S}^{1}}{\max}\,\sum_{j=1}^{m}{|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|}}{2\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}}.$
@@
[21] Suppose $\boldsymbol{u}_{1},\dotsc,\boldsymbol{u}_{m}\in\mathbb{R}^{2}$ are distinct unit vectors that span $\mathbb{R}^{2}$, and suppose that $\alpha_{1},\dotsc,\alpha_{m}>0$. Then there exists a strictly convex $m$-gon P{P} in $\mathbb{R}^{2}$, having edge unit outer normals $\boldsymbol{u}_{1},\dotsc,\boldsymbol{u}_{m}$ and corresponding edge lengths $\alpha_{1},\dotsc,\alpha_{m}>0$, if and only if $\sum_{j=1}^{m}\alpha_{j}\boldsymbol{u}_{j}=\boldsymbol{0}.$ Moreover, such a polygon P{P} is unique up to translation.
@@
Suppose $E=\{\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}\}\subset\mathbb{R}^{2}$. Then there exists a strictly convex $m$-gon P{P} in $\mathbb{R}^{2}$, having the edge set $E$ if and only if $\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}$ are nonzero vectors with pairwise distinct directions and $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}.$ Moreover, such a polygon P{P} is unique up to translation.
@@
Assume that $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$. For all $j\in[m]$, rotating each vector $\frac{\boldsymbol{e}_{j}}{\|\boldsymbol{e}_{j}\|_{2}}$ counterclockwise by $\frac{\pi}{2}$ yields $\boldsymbol{u}_{j}$, while preserving the zero-sum property $\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}\boldsymbol{u}_{j}=\boldsymbol{0}$. It is easy to verify that $\boldsymbol{u}_{1},\dotsc,\boldsymbol{u}_{m}\in\mathbb{R}^{2}$ are distinct unit vectors that span $\mathbb{R}^{2}$. Therefore we can utilize Theorem 2.3 to obtain a strictly convex $m$-gon P{P} with edge unit outer normals $\boldsymbol{u}_{1},\dotsc,\boldsymbol{u}_{m}$ and corresponding edge lengths $\|\boldsymbol{e}_{1}\|_{2},\dotsc,\|\boldsymbol{e}_{m}\|_{2}$, i.e., with edge vectors $\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}$. Moreover, P{P} is unique, disregarding the translation of polygons.
Conversely, assume that P{P} is a strictly convex $m$-gon with the edge set $E$. Assume that $\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}$ are arranged counterclockwise. It is clear that $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$ and each $\boldsymbol{e}_{i}$ is nonzero. Without loss of generality, it is enough to prove that $\boldsymbol{e}_{1}$ and $\boldsymbol{e}_{i}$ have distinct directions for any $i>1$. For each $i\in\{2,\ldots,m\}$, define $\alpha_{i}$ as the counterclockwise angle from $\boldsymbol{e}_{1}$ to $\boldsymbol{e}_{i}$. Since $P$ is strictly convex, we have $\alpha_{i}\in(0,2\pi)$, and $\alpha_{i}$ is strictly increasing as $i$ increases. Thus, there does not exist an integer $i\in\{2,\ldots,m\}$ such that $\alpha_{i}=0$, i.e., $\boldsymbol{e}_{1}$ and $\boldsymbol{e}_{j}$ have the same direction. This completes the proof.